# HLSTester: Efficient Testing of Behavioral Discrepancies with LLMs for High-Level Synthesis

### v1

@@ 3.1. Preprocessing
(1) Code Semantic Analysis: The LLM performs a semantic analysis of the C++ testbench to understand its core logic, functionality, and structure. This step ensures that the LLM fully captures the key functionalities and operations in generating a testbench that is compatible with the HLS tool.
@@
(2) Usage limitations: To meet HLS tool requirements, the LLM adjusts the header files and data types in the original C/C++ testbench to match the definitions provided in the HLS User Guide. For instance, standard data transfers in C++ need to be changed to stream-based data transfers by including the $<$hls_stream.h$>$ header file and using the hls::stream class to emulate hardware stream processing structures. Data types should also be adapted for the HLS program, such as replacing the default int type in the C++ testbench with the ap_int type.
Figure 4. LLM-aided HLS-compatible testbench generation.
@@
To enhance the prompts of LLMs, a Retrieval-Augmented Generation (RAG) technique is employed by incorporating external expert guidance via a retriever. This process involves three stages:
@@
① An external rule library is created containing HLS rule templates. Each template includes potential error logs that may be encountered during synthesis, a brief summary of the associated HLS rule, and usage examples extracted from the HLS tool manual.
@@
② The existing C/C++ testbench is compiled with the HLS tool, and its error logs are used as queries to retrieve matching templates from the rule library. These rule templates embed potential error logs that not only facilitate matching with the compiler-generated error logs but also incorporate guidelines for correcting these errors. A sentence transformer (b17, ) is used to convert the compiler-generated error log and each rule template into embedding vectors. The retrieval process can be formalized as follows:
@@
(2) $E=\mathrm{embed}(\text{error\_log}),\quad R_{i}=\mathrm{embed}(\text{rule\_template}_{i}),\vspace{-0.15cm}$
@@
where $E$ and $R_{i}$ represent the embedding vectors of the compiler-generated error log and the $i$th rule template, respectively. Given an error log with embedding $E$ and a rule template with embedding $R$, their similarity is computed using the cosine similarity, and the rule template with the highest similarity is selected as follows:
@@
(3) $\mathrm{sim}(E,R)=\frac{E\cdot R}{\|E\|\,\|R\|}\quad\text{and}\quad R^{*}=\underset{R_{i}\in\mathcal{R}}{\operatorname{arg\,max}}\,\mathrm{sim}(E,R_{i}),\vspace{-0.18cm}$
@@
③ Once the template with the highest similarity is retrieved, its content is incorporated into the prompt provided to the LLM to generate an HLS-compatible testbench.
@@
To illustrate the application of RAG in generating an HLS compatible testbench, we present an example as shown in Fig. 4. Initially, when the HLS compiler detects incompatible errors (a), such as those related to the use of $<$vector$>$ and dynamic arrays, it generates error logs. These error logs are then used to retrieve the most appropriate rule template from the external library via a sentence transformer (b). Once the rule template with the highest similarity is identified (c), for example, one that recommends replacing $<$vector$>$ with $<$hls_vector.h$>$, it is incorporated into the prompt provided to the LLM (d). The LLM then follows this guideline to generate a new testbench (e). If the HLS compiler still detects an error, the process iteratively repairs the testbench until a correct version that conforms to HLS constraints is obtained.
Figure 5. Using backward slicing technique to pinpoint key variables in the HLS program.
Figure 6. Code instrumentation and spectra monitoring.
3.2. Code Instrumentation and Runtime Spectra Monitoring for Discrepancy Analysis
@@
To enable an in-depth analysis of discrepancies, a backward slicing technique, combined with the semantic understanding capabilities of LLMs, is employed to instrument key variables within the program, and monitoring scripts are developed to record their runtime spectra. Backward slicing is a program analysis technique that traces data dependencies in reverse from a target output to identify all variables that affect its value (b16.5, ). For example, in a function that computes an output through a series of intermediate calculations, backward slicing can identify the relevant variables that contribute to the final result. Consider a dependency graph $G=(V,E)$, where $V$ is the set of variables and $E$ is the set of dependency edges. Let $x$ be the target output variable that needs to be traced back. The backward slicing process is then formalized as follows:
(4) $S_{0}=\{\,x\},\quad S_{i+1}=S_{i}\;\cup\;\bigl{\{}\,u\in V\;\mid\;\exists\,v\in S_{i}:(u,v)\in E\bigr{\}},\vspace{-0.15cm}$
@@
where $S_{0}$ starts with the target output variable $x$, and with each iteration, any variable $u$ that influences a variable $v$ already in the current slice $S_{i}$ is added to form the next slice $S_{i+1}$. The slicing process is repeated until no new variables can be added, i.e., when $S_{i+1}=S_{i}$, and the resulting slice is denoted by $S(x)=S_{i}$. The set of key variables, denoted as $\mathrm{KV}$, is defined as all the variables in the backward slice except the output variable itself:
(5) $\mathrm{KV}(x)=\{\,v\in S(x)\mid v\neq x\},\vspace{-0.15cm}$
meaning any variable that contributes (directly or indirectly) to the final output $x$ is considered a key variable.
To illustrate the process of the backward slicing technique, a simple example is presented in Fig. 5. The process begins at the target output variable x, where discrepancies between C/C++ execution and RTL simulation are observed. Analysis shows that x depends on temp.o1 and c. Since c cannot be traced further, it is identified as a key variable (KV). The slicing then traces temp.o1 in the mul function and reveals its dependence on a, which is subsequently marked as a key variable. The final set of key variables is therefore KV(x) = {temp.o1, c, a}, and these variables are targeted for instrumentation and runtime monitoring.
@@
Figure 7. Test input generation and redundancy-aware filtering.
@@
To facilitate runtime monitoring, our framework leverages an LLM to instrument the program using the previously pinpointed key variables. As shown in Fig. 6, once the key variables are pinpointed, the HLS code is integrated with predefined instrumentation templates, which are provided to the LLM as prompts. These templates specify how to add instrumentation constructs and specify the output format for the runtime spectra. The LLM then generates an instrumented version of the HLS code that records the runtime spectra $Spectra(V)$ of these variables, which can be defined as follows: $Spectra(V)=\{(\mathrm{val}(V),\,\mathrm{offset}(V),\,\mathrm{loop}(V),\,\mathrm{stack}(V),\,\mathrm{FIFO}(V))\}$, where $\mathrm{val}(V)=[\min(V),\,\max(V)]$ denotes the recorded value range, $\text{offset}(V)$ captures array offsets, $\text{loop}(V)$ records loop iteration counts, $\text{stack}(V)$ represents static stack usage, and $\text{FIFO}(V)$ tracks FIFO queue sizes. The runtime spectra are subsequently analyzed by a monitoring script, which generates detailed reports to highlight discrepancy symptoms such as overflow, division by zero, and truncation. Table 1 summarizes the five categories of spectra currently supported by this framework; these mappings can be extended by modifying a configuration file. For instance, array access offsets are monitored to detect out-of-bounds behavior, and actual loop iteration counts are verified to ensure that loop unrolling optimizations align with expectations. Furthermore, the monitoring script reports a feedback array (e.g., spectra_feedback) that records the type, name, minimum, and maximum value of each variable in Table 2. This feedback also provides a basis for subsequent dynamic mutation, which will be detailed in the next subsection.
3.3. Multi-Dimensional Guided Test Input Generation for Effective Testing
@@
To achieve a comprehensive exploration of both breadth and depth in the test input space, two complementary techniques are integrated for test input generation: dynamic mutation and LLM-guided mechanism. As shown in Fig. 7, the dynamic mutation approach broadly explores the test input space by applying a diverse set of mutation operations, while the LLM-guided mechanism uses semantic analysis to generate test inputs that target unique execution paths. Through multiple rounds of validation, 30% of test inputs are generated by the LLM, focusing on program insights, while the remaining test inputs are generated via dynamic mutation to ensure broad coverage.
3.3.1. Dynamic Mutation
@@
As different mutations vary in their ability to trigger different behaviors (b16.6, ; b16.7, ), prioritizing effective mutation operations to test platform-dependent behaviors in HLS presents a significant challenge. In our framework, activation probabilities for mutation types are dynamically adjusted to prioritize those that produce new spectra values, thereby increasing the chances of revealing platform-dependent behavioral discrepancies. Since FPGA test inputs are typically arrays or matrices, our scheme employs eight mutation types to generate new test inputs:
Table 2. Dynamic mutation.
@@
Data Size Mutation: Inserts or deletes random elements in the test input, e.g., changing the input [h, i, j] to [h, i, j, k].
@@
Data Dimension Mutation: Adjusts matrix dimensions, such as adding or removing a row/column, e.g., modifying [[h, i, j], [m, n, p]] to [[h, i, j]] or [[h, i], [m, n]].
@@
Zero Value Mutation: Sets the elements to zero, e.g., changing an element in the array [h, i, j] to 0 to obtain [h, 0, j].
@@
Data Type Mutation: Changes the data type while keeping the same value, e.g., converting the integer 2 to the floating-point 2.0.
Bit Mutation: Randomly flips one or more bits.
Byte Mutation: Randomly flips one or more bytes.
@@
(7) $P_{i}^{(t+1)}=P_{i}^{(t)}-\frac{\alpha}{l-1},\quad i\neq m,\vspace{-0.12cm}$
where $\alpha$ is a predefined update factor set at $0.04$. These updates are subject to the following constraints:
@@
For example, in the second iteration (transitioning from t=1 to t=2), if mutation operation $T_{5}$ generated a test input that expanded the spectra extreme value, its activation probability increases from $P_{5}^{(1)}=0.125$ to $P_{5}^{(2)}=P_{5}^{(1)}+\alpha=0.125+0.04=0.165$, while the other operations adjust to $P_{i}^{(2)}=0.119\quad(i\neq 3)$.
An example illustrating the dynamic mutation guided by spectra feedback is shown in Table 2. The first column lists the test input IDs, while the second column displays the test inputs along with their corresponding mutation types $T_{m}$, where $m\in[1,8]$, and $m\in\mathbb{Z}^{+}$, representing the eight designed mutation types. Columns 3 to 6 record the monitored variable types, names, and the current extreme values captured by the spectra monitoring script, whereas column 7 presents the corresponding real-time spectra ranges of the variables. The final column indicates the activation probability assigned to each mutation type. Initially, a test input seed [1, 1, 7, 5] is employed for HLS testing. In the first mutation round, each mutation type is assigned an activation probability of 0.125. Mutation type $T_{2}$ is activated, yielding a new test input (ID1: [1, 4, 2, 7]). Since the sizes and ranges of the variable spectra remain unchanged, the activation probabilities are maintained. Subsequently, mutation type $T_{5}$ is activated, producing a new test input (ID3: [1, 5, 8, 3]). At this point, the monitoring script detects that the variable sum attains a new extreme value of 17, prompting an increase in the activation probability for mutation type $T_{5}$ in subsequent mutations.
3.3.2. LLM-Guided Test Input Generation
Although dynamic mutation effectively minimizes ineffective mutations while ensuring broad coverage of the test input space, relying solely on mutation may not sufficiently explore the program’s details. At this stage, by leveraging the semantic understanding and adaptive learning capabilities of LLMs, an LLM-guided test input generation mechanism is further proposed. Inspired by the Chain-of-Thought technique, an LLM-driven reasoning chain is employed to progressively analyze discrepancies between an HLS program and its corresponding C/C++ program. As shown in Fig. 8, the process initiates with an overall code analysis to establish an understanding of the program structure. This is followed by a statement-level analysis that delves into the code logic by examining statements and function calls. Finally, a directive analysis phase focuses on hardware-specific interpretations, such as #typedef ap_int<n> and #pragma HLS pipeline, that can customize bit width or alter the execution mode. An example of the LLM-driven reasoning chain, including the associated prompts and responses, is provided at the bottom of Fig. 8, illustrating how the framework analyzes potential discrepancies for the test input generation. The key considerations for employing LLMs to generate test inputs for HLS programs are as follows:
Figure 8. LLM-aided reasoning chain for progressive discrepancy analysis; (a) Execution steps; (b) Analysis example.
(3) Data Format Constraints: HLS programs often have strict data format requirements. The customized data type specifies both the variable type and its bit width (e.g., ap_fixed$<$N$>$ on FPGA), while the data structure defines the format of the input data, such as a one-dimensional array or a two-dimensional matrix, along with its dimensions (e.g., A[Row][Column] on FPGA). The prompts provided to the LLM need to clearly specify these format constraints to ensure that the generated test inputs meet the required specifications. In addition, the prompt should explicitly describe how the test inputs should be arranged. For example, generating one test input per line, separating values with spaces, and placing them between triple backticks. This ensures that the output format is correct and prevents testing failures due to formatting issues.
Table 3. Comparison of the proposed HLSTester with the GPT baseline and the traditional method.
Figure 9. Comparison of the average number of behavioral discrepancies detected within the same time budget across ten real-world tasks, as shown on the x-axes of the figures.
3.4. Redundancy-Aware Filtering of Test Inputs for Efficient Testing
To further reduce hardware testing latency, a redundancy-aware filtering mechanism is proposed to track the value ranges and data sizes of previously tested inputs, thereby avoiding re-simulation of inputs that are likely to exhibit the same behavioral discrepancies.
Figure 10. Comparison of the simulation pass rate of the generated HLS testbench among the proposed HLSTester, GPT baseline, and GPT with RAG.
This mechanism dynamically maintains a global record table that stores the value ranges and data sizes of observed test inputs. When a new test input is generated, the framework first checks whether its value range or size exceeds those currently recorded. If the input’s range or size extends beyond the recorded range, it is likely to trigger new behaviors and hardware simulation will be executed. If the test input remains within the previously recorded range, it implies that the test input will not trigger new behavioral discrepancies, and hardware simulation is skipped. For example, assume that the current recorded input range is (2, 5). If the input is [1, 4, 5], where the minimum value is lower than the recorded range, hardware simulation is executed, and the range is updated to (1, 5). However, if a subsequently generated input is [1, 3, 5], its range remains within (1, 5). Since inputs within the recorded range are likely to produce repetitive known discrepancies, the hardware simulation could be skipped, thereby enhancing testing efficiency by reducing unnecessary simulations.
4. Experimental Results
We demonstrate the results of the proposed HLSTester across 10 real-world tasks in terms of the simulation pass rate of the generated testbench and instrumented program, the number of detected behavioral discrepancies within the same time budget, and the total time required for testing all behavioral discrepancies. Our benchmarks are from related work (b16.1, ; b16.2, ; b16.3, ) covering different hardware directives (e.g., customized bit width, memory management, parallel operation, etc.), thereby exhibiting diverse behavior discrepancies.
[…]
@@ 4. Experimental Results
Figure 11. Comparison of the simulation pass rate of the instrumented HLS code among the proposed method, GPT baseline, and GPT with backward slicing.
@@
Figure 12. Comparison of the pass rate for instrumented C++ code between the proposed method and the GPT baseline.
@@
Table 3 compares the performance of the proposed HLSTester with the GPT baseline and the traditional testing method (b8, ). The first column lists the name of benchmark tasks, while the second through seventh columns show the simulation pass rates of the generated testbench and instrumented code for HLS, with the proposed framework demonstrating superior performance compared to the baseline. The eighth through tenth columns report the average number of detected behavioral discrepancies, demonstrating that the proposed HLSTester identifies more discrepancies within the same time budget. The final four columns detail the total time required to test all behavioral discrepancies and the corresponding reductions. According to these columns, the proposed HLSTester outperforms both the GPT baseline and the traditional method by achieving a higher pass rate and shorter overall testing time.
@@
To demonstrate the effectiveness of the proposed method in testbench generation for HLS, we compare the simulation pass rate of the proposed HLSTester with using GPT directly. According to Fig. 10, the proposed HLSTester outperforms the baseline by achieving an average 20.67% increase in testbench generation.
@@
Fig. 11 and Fig. 12 present comparisons of the simulation pass rate for code instrumentation, where the proposed framework was compared with the direct use of GPT and GPT with backward slicing. In these comparisons, the proposed LLM-aided framework demonstrated average improvements of 25.33% and 18.67% in HLS and C/C++ code instrumentation, respectively.
@@
To demonstrate the acceleration achieved by the HLSTester in generating effective test inputs for efficient behavioral discrepancy detection, four baselines, including ablation options, were created by downgrading HLSTester in Fig. 9, which are summarized as follows:

### v3

@@ III-A Testbench Modification for HLS
(1) Code Semantic Analysis: The LLM performs a semantic analysis of the C++ testbench to understand its core logic, functionality, and structure. This step ensures that the LLM fully captures the key functionalities and operations in modifying a testbench that is compatible with the HLS tool.
@@
(2) Usage limitations: To meet HLS tool requirements, the LLM adjusts the header files and data types in the original C/C++ testbench to match the definitions provided in the HLS User Guide. For instance, standard data transfers in C++ need to be changed to stream-based data transfers by including the $<$hls_stream.h$>$ header file and using the hls::stream to emulate hardware stream processing structures. Data types should also be adapted for the HLS program, such as replacing the int type in the C++ testbench with the ap_int type.
Figure 4: LLM-aided testbench modification for HLS.
@@
To enhance LLM prompts, a Retrieval-Augmented Generation (RAG) technique is employed by incorporating external guidance via a retriever. This process involves three stages:
@@
First, an external rule library is created containing HLS rule templates. Each template includes potential error logs that may be encountered during synthesis, a summary of the associated HLS rule, and usage examples extracted from the HLS manual.
@@
Second, the existing C/C++ testbench is compiled with the HLS tool, and its error logs are used as queries to retrieve templates from the rule library. These rule templates embed potential error logs that not only facilitate matching with the compiler-generated error logs but also incorporate guidelines for correcting these errors. A sentence transformer [39] is used to convert the error log and each rule template into embedding vectors. The retrieval process can be formalized as follows:
@@
$E=\mathrm{embed}(\text{error\_log}),\quad R_{i}=\mathrm{embed}(\text{rule\_template}_{i}),\vspace{-0.15cm}$ (1)
@@
where $E$ and $R_{i}$ represent the embedding vectors of the compiler-generated error log and the $i$th rule template. Given an error log embedding $E$ and a rule template embedding $R$, their similarity is computed via cosine similarity, and the rule template with the highest similarity is selected as follows:
@@
$\mathrm{sim}(E,R)=\frac{E\cdot R}{\|E\|\,\|R\|}~{},\quad R^{*}=\underset{R_{i}\in\mathcal{R}}{\operatorname{arg\,max}}\,\mathrm{sim}(E,R_{i}),\vspace{-0.1cm}$ (2)
@@
Third, once the template with the highest similarity is retrieved, it is incorporated into the LLM prompt to guide the modification of the testbench into an HLS-compatible version.
@@
An example of applying RAG to modify a C/C++ testbench into an HLS-compatible version is shown in Fig. 4. Initially, when the HLS compiler detects incompatible errors (a), such as those related to the use of $<$vector$>$ and dynamic arrays, it generates error logs. These error logs are then used to retrieve the most appropriate rule template from the external library via a sentence transformer (b). Once the rule template with the highest similarity is retrieved (c), for example, one that recommends replacing $<$vector$>$ with $<$hls_vector.h$>$, it is incorporated into the prompt provided to the LLM (d). The LLM then follows this guideline to generate a new testbench (e). If the HLS compiler still detects an error, the process iteratively repairs the testbench until a correct version that conforms to HLS constraints is obtained.
Figure 5: Using backward slicing to pinpoint key variables in the HLS program.
Figure 6: Code instrumentation and spectra monitoring for discrepancy analysis.
III-B Instrumentation and Monitoring for Discrepancy Analysis
@@
To enable an in-depth analysis of discrepancies, a backward slicing technique, combined with the semantic understanding capabilities of LLMs, is employed to instrument key variables within the program, and monitoring scripts are developed to record their runtime spectra. Backward slicing is a program analysis technique that traces data dependencies in reverse from a target output to identify all variables that affect its value [36]. For example, in a function that computes an output through a series of intermediate calculations, backward slicing can identify the relevant variables that contribute to the final result. Consider a dependency graph $G=(V,E)$, where $V$ is the set of variables and $E$ is the set of dependency edges. Let $x$ be the target output variable that needs to be traced back. The backward slicing process is then formalized as follows:
$S_{0}=\{\,x\},S_{i+1}=S_{i}\cup\bigl{\{}\,u\in V\mid\exists\,v\in S_{i}:(u,v)\in E\bigr{\}},\vspace{-0.16cm}$ (3)
@@
where $S_{0}$ starts with the target output variable $x$, and with each iteration, any variable $u$ that influences a variable $v$ already in the current slice $S_{i}$ is added to form the next slice $S_{i+1}$. The slicing process is repeated until no variables can be added, i.e., when $S_{i+1}=S_{i}$, and the final slice is denoted by $S(x)=S_{i}$.
@@
Figure 7: Multi-dimensional test input generation and redundancy-aware filtering.
To illustrate the process of the backward slicing technique, a simple example is presented in Fig. 5. The process begins at the target output variable x, where discrepancies between C/C++ execution and RTL simulation are observed. Analysis shows that x depends on temp.o1 and c. Since c cannot be traced further, it is identified as a key variable (KV). The slicing then traces temp.o1 in the mul function and reveals its dependence on a, which is subsequently marked as a key variable. The final set of key variables is KV(x) = {temp.o1, c, a}, and these variables are targeted for instrumentation and runtime monitoring.
@@
To facilitate runtime monitoring, our framework leverages an LLM to instrument the program using the previously pinpointed key variables. As shown in Fig. 6, once the key variables are pinpointed, the HLS code is integrated with predefined instrumentation templates, which are provided to the LLM as prompts. These templates specify how to add instrumentation constructs and specify the output format for the runtime spectra. The LLM then generates an instrumented version of the HLS code that records the runtime spectra $Spectra(V)$ of these variables, which can be defined as follows: $Spectra(V)=\{(\mathrm{val}(V),\,\mathrm{offset}(V),\,\mathrm{loop}(V),\,\mathrm{stack}(V),\,\mathrm{FIFO}(V))\}$, where $\mathrm{val}(V)=[\min(V),\,\max(V)]$ denotes the recorded value range, $\text{offset}(V)$ captures array offsets, $\text{loop}(V)$ records loop iteration counts, $\text{stack}(V)$ represents static stack usage, and $\text{FIFO}(V)$ tracks FIFO queue sizes. The runtime spectra are subsequently analyzed by a monitoring script, which generates detailed reports to highlight discrepancy symptoms such as overflow, division by zero, and truncation. Table I summarizes the five categories of spectra currently supported by this framework; these mappings can be extended by modifying a configuration file. For instance, array access offsets are monitored to detect out-of-bounds behavior, and actual loop iteration counts are verified to ensure that loop unrolling optimizations align with expectations. Furthermore, the monitoring script reports a feedback array (e.g., spectra_feedback) that records the type, name, minimum, and maximum value of each variable in Table III-C1. This feedback also provides a basis for subsequent dynamic mutation, which is detailed in the next subsection.
III-C Multi-Dimensional Guided Test Input for Effective Testing
@@
At this stage, two complementary techniques are integrated for test input generation: dynamic mutation and LLM-guided mechanism. As shown in Fig. 7, the dynamic mutation approach broadly explores the test input space by applying a diverse set of mutation operations, while the LLM-guided mechanism uses semantic analysis to generate test inputs that target program features. Through multiple rounds of validation, 30% of test inputs are generated by the LLM, focusing on program insights, while the remaining test inputs are generated via dynamic mutation to ensure broad coverage.
III-C1 Dynamic Mutation
@@
As different mutations vary in their ability to trigger different behaviors [37], prioritizing effective mutation operations to test platform-dependent behaviors in HLS presents a significant challenge. In our framework, activation probabilities for mutation types are dynamically adjusted to prioritize those that produce new spectra values, thereby increasing the chances of revealing platform-dependent behavioral discrepancies. Since FPGA test inputs are typically arrays or matrices, our scheme employs eight mutation types to generate new test inputs:
TABLE II: DYNAMIC MUTATION FOR TEST INPUT GENERATION.
@@
Data Size Mutation: Inserts or deletes elements in the test input, e.g., changing the input [i, j] to [i, j, k].
@@
Data Dimension Mutation: Adjusts matrix dimensions, such as adding or removing a row/column, e.g., modifying [[i, j], [m, n]] to [[i, j]] or [[i], [m]].
@@
Zero Value Mutation: Sets the elements to zero, e.g., changing an element of [i, j] to 0 to obtain [0, j].
@@
Data Type Mutation: Changes the data type while keeping the same value, e.g., converting integer 2 to floating-point 2.0.
Bit Mutation: Randomly flips a bit, e.g., 0101 to 0100.
Byte Mutation: Randomly flips a byte, e.g., turning the 16‑bit value 0x1A2B into 0x1AD4 by flipping its lower byte.
@@
$P_{i}^{(t+1)}=P_{i}^{(t)}-\frac{\alpha}{l-1},\quad i\neq m,\vspace{-0.18cm}$ (5)
where $\alpha$ is a predefined update factor set at $0.04$. For example, in the second iteration (transitioning from t=1 to t=2), if mutation operation $T_{5}$ generated a test input that expanded the spectra extreme value, its activation probability increases from $P_{5}^{(1)}=0.125$ to $P_{5}^{(2)}=P_{5}^{(1)}+\alpha=0.125+0.04=0.165$, while the other operations adjust to $P_{i}^{(2)}=0.119\quad(i\neq 3)$.
An example illustrating the dynamic mutation guided by spectra feedback is shown in Table III-C1. The first column lists the test input IDs, while the second column displays the test inputs along with their corresponding mutation types $T_{m}$, where $m\in[1,8]$, and $m\in\mathbb{Z}^{+}$, representing the eight designed mutation types. Columns 3 to 6 record the monitored variable types, names, and the current extreme values captured by the spectra monitoring script, whereas column 7 presents the real-time spectra ranges of the variables. The final column indicates the activation probability assigned to each mutation. Initially, a test input seed [1, 1, 7, 5] is employed for HLS testing. In the first mutation round, each mutation type is assigned an activation probability of 0.125. Mutation type $T_{5}$ is activated, yielding a new test input (ID1: [1, 4, 7, 2]). Since the ranges of the variable spectra have not been expanded, the activation probabilities are maintained. Subsequently, mutation type $T_{7}$ is activated, producing a new test input (ID3: [1, 4, 7, 8]). At this point, the script detects that the variable sum attains a new extreme value of 20, prompting an increase in the activation probability for mutation type $T_{7}$ in subsequent mutations.
III-C2 LLM-Guided Test Input Generation
Although dynamic mutation effectively minimizes ineffective mutations while ensuring broad coverage of the test input space, relying solely on mutation may not sufficiently explore the program’s details. At this stage, by leveraging the semantic understanding and adaptive learning capabilities of LLMs, an LLM-guided test input generation mechanism is further proposed. Inspired by the Chain-of-Thought technique, an LLM-driven reasoning chain is employed to progressively analyze discrepancies between an HLS program and its corresponding C/C++ program. As shown in Fig. 8, the process initiates with an overall code analysis to establish an understanding of the program structure. This is followed by a statement-level analysis that delves into the code logic by examining statements and function calls. Finally, a directive analysis phase focuses on hardware-specific interpretations, such as #typedef ap_int<n> and #pragma HLS pipeline, that can customize bit width or alter the execution mode. An example of the LLM-driven reasoning chain, including the prompts and responses, is provided at the bottom of Fig. 8, illustrating how the framework analyzes potential discrepancies for the test input generation. The considerations for employing LLMs to generate test inputs are as follows:
Figure 8: LLM-aided reasoning chain for progressive behavioral discrepancy analysis. Top: Execution steps. Bottom: Analysis example.
(3) Data Format Constraints: HLS programs often have strict data format requirements. The customized data type specifies both the variable type and its bit width [38] (e.g., ap_fixed$<$N$>$ on FPGA), while the data structure defines the format of the input data, such as a one-dimensional array or a two-dimensional matrix, along with its dimensions (e.g., A[Row][Column] on FPGA). The prompts provided to the LLM need to clearly specify these format constraints to ensure that the generated test inputs meet the required specifications. In addition, the prompt should explicitly describe how the test inputs should be arranged. For example, generating one test input per line, separating values with spaces, and placing them between triple backticks. This ensures that the output format is correct and prevents testing failures due to formatting issues.
TABLE III: COMPARISON OF THE PROPOSED HLSTESTER WITH THE GPT BASELINE AND THE TRADITIONAL METHOD.
Figure 9: Comparison of the number of behavioral discrepancies detected within the same time budget across ten tasks, as shown on the x-axes of the figures.
III-D Redundancy Filtering of Test Inputs for Efficient Testing
To avoid redundant simulations of test inputs that would exhibit the same behavioral discrepancies, a redundancy-aware technique is proposed to track the value ranges and data sizes of previously tested inputs. This technique dynamically maintains a global record table that stores the value ranges and data sizes of observed test inputs. When a new test input is generated, the framework first checks whether its value range or size exceeds those currently recorded. If the input’s range or size extends beyond the recorded range, it is likely to trigger new behaviors and hardware simulation will be executed. If the test input remains within the previously recorded range, it implies that the test input will not trigger new discrepancies, and simulation is skipped. For example, assume that the current recorded input range is (2, 5). If the input is [1, 4, 5], where the minimum value is lower than the recorded range, simulation is executed, and the range is updated to (1, 5). However, if a subsequently generated input is [1, 3, 5], its range remains within (1, 5). Since inputs within the recorded range are likely to produce repetitive discrepancies, the hardware simulation could be skipped, thereby reducing unnecessary simulations.
IV Experimental Results
We demonstrate the results of the proposed LLM-aided framework for HLS across 10 tasks in terms of the simulation pass rate of the modified testbench and instrumented code, the number of detected behavioral discrepancies within the same time budget, and the total time required for testing all behavioral discrepancies. Our benchmarks are from related work [31, 32, 33, 34] covering different hardware directives (e.g., customized bit width, memory management, parallelism, etc.).
Figure 10: Comparison of the simulation pass rate of the modified HLS testbench among the proposed HLSTester, GPT baseline, and GPT with RAG.
@@ IV Experimental Results
Figure 11: Comparison of the simulation pass rate of the instrumented HLS code among the proposed method, GPT baseline, and GPT with backward slicing.
@@
Figure 12: Comparison of the pass rate for instrumented C++ code between the proposed method and the GPT baseline.
@@
Table III-C2 compares the performance of the proposed HLSTester with the GPT baseline and the traditional testing method [12]. The first column lists the name of benchmark tasks, while the second through seventh columns show the simulation pass rates of the modified testbench and instrumented code for HLS, with the proposed framework demonstrating superior performance compared to the baseline. The eighth through tenth columns report the average number of detected behavioral discrepancies, demonstrating that the proposed HLSTester identifies more discrepancies within the same time budget. The final four columns detail the total time required to test all discrepancies and the corresponding reductions. According to these columns, the proposed HLSTester outperforms both the GPT baseline and the traditional method by achieving a higher pass rate and shorter overall testing time.
@@
To demonstrate the effectiveness of the proposed method in testbench modification, we compare the simulation pass rate of the proposed HLSTester with using GPT directly. According to Fig. 10, the HLSTester outperforms the baseline by achieving an average 20.67% increase in testbench modification.
@@
Fig. 11 and Fig. 12 present comparisons of the simulation pass rate for code instrumentation, where the proposed framework was compared with the direct use of GPT and GPT with backward slicing. In these comparisons, the proposed HLSTester demonstrated average improvements of 25.33% and 18.67% in HLS and C/C++ code instrumentation, respectively.
@@
To demonstrate the acceleration achieved by the HLSTester in generating effective test inputs for efficient discrepancy detection, four baselines including ablation options were created by downgrading HLSTester in Fig. 9, summarized as follows: